OpenStudy (anonymous):

Given the ordered pair (x,2) what value of x would make the ordered pair a solution of the equation 5x-3y=19?

OpenStudy (anonymous):

plug in 2 for y and solve

OpenStudy (anonymous):

I really dont know how to do it at all

OpenStudy (anonymous):

ok lets start out by solving the equation for y

OpenStudy (anonymous):

Ok how do you do that?

OpenStudy (anonymous):

start by getting y by itself

OpenStudy (anonymous):

Ya ok.. well how do you do that?

OpenStudy (anonymous):

ok instead of going through all of this i will show you the steps

OpenStudy (anonymous):

Ok.(:

OpenStudy (anonymous):

5x-3y=19 5x - 3(2) = 19 5x - 6 = 19 5x = 25 x = 5

OpenStudy (anonymous):

make sense?

OpenStudy (anonymous):

Yeah thank you.(:

OpenStudy (anonymous):

not a problem